# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ 6.  Conclusion and Open Problems
We studied bidding in repeated uniform price auctions for a value maximizing buyer with per-round RoI constraints and characterized safe bidding strategies that can be efficiently learnt and are robust against various strong benchmarks. This study suggests several interesting future research directions: (1) analyzing bidding strategies in discriminatory-price (“pay-as-bid”) auctions; (2) modeling bidders with time-varying valuations and cumulative RoI constraints; (3) incorporating both RoI and budget constraints for a more realistic model of bidder behavior; and (4) extending the analysis to the more complex setting of combinatorial auctions with non-identical items.
@@ Appendix A Omitted Proofs from Section 3
Overbidding is not a safe bidding strategy. To see this, let $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell}),\dots,(b_{m},q_{m})\rangle$ be an overbid such that $b_{\ell}>w_{Q_{\ell}}$. Now, consider an auction in which the competing bids are:
@@
We show that for any $m\in\mathbb{N}$, most safe bidding strategies are very weakly dominated for which they can removed from $\mathscr{S}_{m}$, resulting in $\mathscr{S}_{m}^{\star}$. We begin by establishing a general result regarding the monotonocity of feasible bid vectors (not necessarily $m$-uniform strategies) for value maximizing bidders. As the result holds for any bid vector, it is also true for $m$-uniform bidding strategies.
@@ Appendix B Omitted Proofs From Section 4
B.1.  Proof of Lemma 4.1
By Theorem 3.2, a strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle\in\mathscr{U}_{m}^{\star}$ for some $\ell\in[m]$ can be uniquely identified by $\{Q_{1},\dots,Q_{\ell}\}$ where $Q_{j}=\sum_{k\leq j}q_{k}$ because $b_{j}=w_{Q_{j}}$ for all $j\in[\ell]$. Let the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. Then, for any round $t$,
$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{j=1}^{\ell}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}$
Hence, the problem (Offline) can be equivalently expressed as
B.2.  Proof of Theorem 4.2
$\mathcal{G}(V,E)$ is a DAG. For convenience, let $d=(m+1,0)$. In the constructed graph, any node $(\ell,j)$ where $\ell\in[m]$ has edges edges either to the next layer, i.e. nodes of the form $(\ell+1,j^{\prime})$ where $j^{\prime}>j$ or to the destination node $(m+1,0)$. Hence, the directed graph has a topological sorting of the nodes implying that $\mathcal{G}(V,E)$ is a DAG.
@@
$\mathsf{w}(\mathfrak{p})$ $=\sum_{e\in\mathfrak{p}}\mathsf{w}(e)=\sum_{\ell=1}^{m}\mathsf{w}((\ell-1,z_{\ell-1})\to(\ell,z_{\ell}))$
=(4)${{=}}\sum_{\ell=1}^{m}\sum_{t=1}^{T}\sum_{k=z_{\ell-1}+1}^{z_{\ell}}v_{k}\cdot\mathbb{I}{\left[w_{z_{\ell}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}$
@@
B.3.  Proof of Theorem 4.3
We begin by observing that there is a bijective mapping between $s$-$d$ paths in the DAG $\mathcal{G}^{t}(V,E)$ and bidding strategies $\mathbf{b}\in\mathscr{U}_{m}^{\star}$ (see proof of Theorem 4.2). The core idea of the proof is to show that Algorithm 1 is an equivalent and efficient implementation of the Hedge algorithm (Freund and Schapire, 1997) where every $s$-$d$ path is treated as an individual expert. To this end, we use an approach proposed in Takimoto and Warmuth (2003) and later used by Brânzei et al. (2023); Potfer et al. (2024); Galgana and Golrezaei (2024). For the sake of completeness, we provide the details.
In Algorithm 1, $\varphi^{t}(u\to\cdot)$ denotes the probability distribution over the outgoing neighbors of node $u$. By the recursive sampling of nodes (Line 4-7 of Algorithm 1), we get that the probability of selecting $s$-$d$ path $\mathfrak{p}$ is
@@
$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)},$ (11)
where $\Gamma^{t-1}(d)=1$ and $\Gamma^{t-1}(\cdot)$ is computed recursively in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$ as follows:
$\Gamma^{t-1}(u)=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))\,.$ (12)
Now, consider a naïve implementation of the Hedge algorithm with learning rate $\eta$ in which each $s$-$d$ path is treated as an individual expert. For $t=1$, $\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)$. For $t\geq 2$, the probability of selecting path $\mathfrak{p}$ in round $t$ is
$\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})=\frac{\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\exp(\eta\sum_{e\in\mathfrak{p}}\mathsf{w}^{t-1}(e))}{\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\exp(\eta\sum_{e\in\mathfrak{p}^{\prime}}\mathsf{w}^{t-1}(e))}\,.$
Now, we show that for any path $\mathfrak{p}$ and $t\in[T]$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. To this end, we first present the following result:
@@
$\Gamma^{t-1}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))$
@@
We prove the result by induction in a bottom-to-top order. For the base case, $u=d$, $\Gamma^{t-1}(d)=1$. Suppose the result is true for all the nodes in layer $\ell+1$ for some $0\leq\ell\leq m$. By the recursion in Eq. 12, for any node $u$ in layer $\ell$,
$\Gamma^{t-1}(u)$ $=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))$
$=\sum_{v:u\to v\in E}\Big{(}\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\Big{)}\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))$
$=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e)),$
@@
We will show by induction on $t\in[T]$ that for any path $\mathfrak{p}$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. For $t=1$ and any path $\mathfrak{p}$, $\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\mathbb{P}^{1}(\mathfrak{p})$. Here, the second equality holds by Eq. 10. Suppose the result holds for round $t-1$, i.e., $\mathbb{P}^{t-1}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)$. For round $t$,
@@
$\Gamma^{t-1}(s)=\sum_{\mathfrak{p}^{\prime}}\prod_{e\in\mathfrak{p}^{\prime}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))=\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\cdot\exp\Big{(}\eta\sum_{e\in\mathfrak{p}^{\prime}}\mathsf{w}^{t-1}(e)\Big{)}$
@@
Substituting the values gives $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$ implying Algorithm 1 is a correct implementation of the Hedge algorithm.
B.3.1.  Bandit Feedback Model
In the bandit setting, the bidder learns about their own allocation only for which $\mathsf{w}^{t}(e)$ can not be exactly computed for all the edges. Hence, we use an unbiased estimator, $\widehat{\mathsf{w}}^{t}(e)$, of $\mathsf{w}^{t}(e)$ as follows:
$\widehat{\mathsf{w}}^{t}(e)$ $=\overline{\mathsf{w}}(e)-\frac{\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}\right]}\quad\text{where}\quad p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\,.$
Here, $p^{t}(e)$ is the probability of selecting edge $e$ in round $t$, $\mathbb{P}^{t}$ is the distribution over all $s$-$d$ paths in $\mathcal{G}^{t}(V,E)$ and $\overline{\mathsf{w}}(e)$ is an edge-dependent upper bound of the edge weight $\mathsf{w}^{t}(e)$. Formally,
$\overline{\mathsf{w}}(e)=\begin{cases}j^{\prime}-j,&\text{if }e=(\ell,j)\to(\ell+1,j^{\prime})\\ 0,&\text{if }e=(\ell,j)\to d\,.\end{cases}$
The algorithm followed by the bidder in the bandit setting is the same as Algorithm 1, except in Line 9, the bidder sets the edge weights as $\widehat{\mathsf{w}}^{t}(e)$ instead of $\mathsf{w}^{t}(e)$. Clearly, $\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)]=\mathsf{w}^{t}(e)$. The unbiased estimator $\widehat{\mathsf{w}}^{t}(e)$ is different from the natural importance-weighted estimator $\frac{\mathsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}\right]}$ for technical reasons (similar to Lattimore and Szepesvári (2020, Equation 11.6)) which we will clarify shortly.
B.3.2.  Regret Upper Bound
@@
Full Information Setting. Having shown that Algorithm 1 emulates the Hedge algorithm, we now recall the its regret bound from Brânzei et al. (2023, Corollary 1) which states that for learning rate $\eta$:
$\mathsf{REG}=-\frac{1}{\eta}\max_{\mathfrak{p}}\log(\mathbb{P}^{1}(\mathfrak{p}))+\frac{\eta TM^{2}}{8}\,.$
Observe that $\mathbb{P}^{1}(\mathfrak{p})\geq M^{-m}$, which implies,
$\mathsf{REG}\leq\frac{m\log M}{\eta}+\frac{\eta TM^{2}}{8}\,.$
Setting $\eta=\frac{1}{M}\sqrt{\frac{8m\log M}{T}}$, we get $\mathsf{REG}\leq O(M\sqrt{mT\log M})$.
Bandit Setting. Using the standard analysis for EXP3 algorithm (see Lattimore and Szepesvári (2020, Chapter 11)), we get
@@
$\mathsf{REG}=-\frac{1}{\eta}\max_{\mathfrak{p}}\log(\mathbb{P}^{1}(\mathfrak{p}))+\sum_{t=1}^{T}\mathbb{E}\left[\frac{1}{\eta}\log\left(\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})e^{\eta\widehat{\mathsf{w}}^{t}(\mathfrak{p})}\right)-\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\mathsf{w}}^{t}(\mathfrak{p})\right]$
where $\widehat{\mathsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\mathsf{w}}^{t}(e)$. Note that $\widehat{\mathsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\mathsf{w}}^{t}(e)\leq\sum_{e\in\mathfrak{p}}\overline{\mathsf{w}}(e)\leq M$ and $\mathbb{P}^{1}(\mathfrak{p})\geq M^{-m}$. For $\eta\leq\frac{1}{M}$,
$\mathsf{REG}$ $\leq\frac{m\log M}{\eta}+\eta\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\mathbb{E}[\widehat{\mathsf{w}}^{t}(\mathfrak{p})^{2}]$
$\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$
@@
where in the first inequality, we used $e^{x}\leq 1+x+x^{2}$ for $x\leq 1$ and $\log(1+x)\leq x$ for $x\geq 0$ and the second inequality follows from Cauchy-Schwarz. Observe that
$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]=\sum_{e}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{e}p^{t}(e)\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}],$
@@
$\sum_{e}p^{t}(e)\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$ $=\sum_{e}p^{t}(e)\Big{[}\overline{\mathsf{w}}(e)^{2}(1-p^{t}(e))+\Big{(}\overline{\mathsf{w}}(e)-\frac{\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e)}{p^{t}(e)}\Big{)}^{2}p^{t}(e)\Big{]}$
$=\sum_{e}\mathsf{w}^{t}(e)^{2}p^{t}(e)+(\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e))^{2}(1-p^{t}(e))\leq\sum_{e}\overline{\mathsf{w}}(e)^{2}\,.$
We now aim to upper bound $\sum_{e}\overline{\mathsf{w}}(e)^{2}$.
$\sum_{e}\overline{\mathsf{w}}(e)^{2}=\sum_{\ell=0}^{m-1}\sum_{j<j^{\prime}\leq M}\overline{\mathsf{w}}((\ell,j)\to(\ell+1,j^{\prime}))^{2}=m\sum_{j<j^{\prime}\leq M}(j-j^{\prime})^{2}\leq 4m\sum_{j=1}^{M}j^{2}\leq 4mM^{3}\,.$
Hence,

### v4

@@ 7. Conclusion and Open Problems
In this work, we studied the bidding problem in repeated uniform-price auctions for a value-maximizing buyer subject to per-round return-on-investment (RoI) constraints. To address this, we introduced the notion of safe bidding strategies as those that satisfy RoI constraints regardless of competing bids and characterized the class of $m$-uniform safe strategies. These strategies depend solely on the buyer’s valuation curve, and the buyer can focus on a finite subset of this class without loss of generality. We proposed a polynomial-time online learning algorithm to efficiently learn optimal strategies within this class. We then analyzed the robustness of safe strategies and the proposed algorithm against stronger benchmarks, showing that our approach achieves $\alpha$-approximate sublinear regret, where $\alpha$ is the richness ratio. We complemented the theoretical results by simulations on semi-synthetic EU ETS auction data which demonstrate that empirical richness ratios are substantially better than the worst-case bounds. We also showed the flexibility and adaptability of safe strategies and our learning algorithm across several model extensions such as adaptive adversaries, sliding-window RoI constraints, and time-varying valuations highlighting their practical relevance.
From a practitioner’s perspective, safe bidding strategies are intuitive, robust, and adaptable to a wide range of risk preferences. The proposed algorithm is simple to implement and backed by strong theoretical guarantees, making it well-suited for real-world deployment. This study opens up several exciting avenues for future investigation. A natural extension is to characterize safe strategies for value-maximizing buyers in other multi-unit auction formats, such as discriminatory-price (pay-as-bid) auctions. From a modeling standpoint, incorporating both RoI and budget constraints could lead to richer and more realistic representations of bidder behavior. While our analysis focuses on a single buyer, it would be valuable to explore market dynamics, equilibrium outcomes, and convergence properties when all participants adopt safe strategies—paralleling recent developments under different market conditions and bidders’ behavior (Conitzer et al., 2022; Li and Tang, 2024; Lucier et al., 2024; Fikioris and Tardos, 2023; Paes Leme et al., 2024). Finally, extending this framework to combinatorial auctions with heterogeneous items presents a host of new algorithmic and design challenges.
Acknowledgements
N.G. and S.S. are partially supported by the MIT Junior Faculty Research Assistance Grant and the Young Investigator Award from the Office of Naval Research (ONR), Award No. N00014-21-1-2776.
@@ Appendix A Omitted Proofs from Section 3
Overbidding is not a safe bidding strategy. To see this, let
$\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell}),\dots,(b_{m},q_{m})\rangle$
be an overbid such that $b_{\ell}>w_{Q_{\ell}}$. Now, consider an auction in which the competing bids are:
@@
We show that for any $m\in\mathbb{N}$, most safe bidding strategies are very weakly dominated for which they can eliminated from $\mathscr{S}_{m}$, resulting in $\mathscr{S}_{m}^{\star}$. We begin by establishing a general result regarding the monotonocity of feasible bid vectors (not necessarily $m$-uniform strategies) for value maximizing bidders. As the result holds for any bid vector, it is also true for $m$-uniform bidding strategies.
@@ Appendix B Omitted Proofs From Section 4
B.1. Proof of Theorem 4.2
$\mathcal{G}(V,E)$ is a DAG. For convenience, let $d=(m+1,\infty)$. In the constructed graph, any node $(\ell,j)$ where $\ell\in[m]$ has edges edges either to the next layer, i.e. nodes of the form $(\ell+1,j^{\prime})$ where $j^{\prime}>j$ or to the destination node $(m+1,\infty)$. Hence, the directed graph has a topological sorting of the nodes implying that $\mathcal{G}(V,E)$ is a DAG.
@@
$\textsf{w}(\mathfrak{p})$ $=\sum_{e\in\mathfrak{p}}\textsf{w}(e)=\sum_{\ell=1}^{m}\textsf{w}((\ell-1,z_{\ell-1})\to(\ell,z_{\ell}))$
[…]
@@
B.2. Proof of Theorem 4.3
We begin by observing that there is a bijective mapping between $s$-$d$ paths in the DAG $\mathcal{G}^{t}(V,E)$ and bidding strategies $\mathbf{b}\in\mathscr{U}_{m}^{\star}$ (see proof of Theorem 4.2). The core idea of the proof is to show that Algorithm 1 is an equivalent and efficient implementation of the Hedge algorithm (Freund and Schapire, 1997) where every $s$-$d$ path is treated as an expert.
In Algorithm 1, $\varphi^{t}(u\to\cdot)$ denotes the probability distribution over the outgoing neighbors of node $u$. By the recursive sampling of nodes, we get that the probability of selecting $s$-$d$ path $\mathfrak{p}$ is
@@
$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)},$ (30)
where $\Gamma^{t-1}(d)=1$ and $\Gamma^{t-1}(\cdot)$ is computed recursively in bottom-to-top fashion for every node $u\in V$ as follows:
$\Gamma^{t-1}(u)=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\textsf{w}^{t-1}(u\to v))\,.$ (31)
Now, consider a naïve version of the Hedge algorithm with learning rate $\eta$ in which each $s$-$d$ path is treated as an individual expert. For $t=1$, the algorithm samples a path uniformly at random from $\mathscr{P}$, the collection of all $s$-$d$ paths in the DAG. For $t\geq 2$, the probability of selecting path $\mathfrak{p}$ in round $t$ is
$\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})=\frac{\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\exp(\eta\sum_{e\in\mathfrak{p}}\textsf{w}^{t-1}(e))}{\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\exp(\eta\sum_{e\in\mathfrak{p}^{\prime}}\textsf{w}^{t-1}(e))}\,.$
We prove Algorithm 1 is equivalent to the Hedge algorithm by showing that $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p}),\forall\mathfrak{p}\in\mathscr{P},\forall t\in[T]$. To this end, we first present the following result:
@@
$\Gamma^{t-1}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))$
@@
We prove the result by strong induction in a bottom-to-top order. For the base case, $u=d$, $\Gamma^{t-1}(d)=1$. Suppose the result is true for all the nodes in layer $\ell^{\prime}$ for some $\ell<\ell^{\prime}$. By the recursion in Eq. 31, for any node $u$ in layer $\ell$,
$\Gamma^{t-1}(u)$ $=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\textsf{w}^{t-1}(u\to v))$
$=\sum_{v:u\to v\in E}\Big(\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\Big)\varphi^{t-1}(u\to v)\cdot\exp(\eta\textsf{w}^{t-1}(u\to v))$
$=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e)),$
@@
Claim B.2.
$\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\frac{1}{|\mathscr{P}|}=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\mathbb{P}^{1}(\mathfrak{p})$.
Proof.
The first equality holds from the definition of the naïve Hedge algorithm above and the last equality follows from Eq. 29. To complete the proof, we need to show that $\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\frac{1}{|\mathscr{P}|}$. As $\varphi^{0}(e)=1$ and $\textsf{w}^{0}(e)=0$ for all $e\in E$ by initialization in Algorithm 1, for any edge $e=u\to v$, we get that
@@
for any $s$-$d$ path $\mathfrak{p}\in\mathscr{P}$. For any node $u\in V$, by B.1, $\Gamma^{0}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}1=|\mathscr{P}(u)|$, the number of paths that start at $u$ and terminate at $d$. By initialization, $\Gamma^{0}(d)=1$ and $\Gamma^{0}(s)=|\mathscr{P}(s)|=|\mathscr{P}|$, which proves the result. ∎
We will show by induction on $t\in[T]$ that for any path $\mathfrak{p}\in\mathscr{P}$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. For $t=1$ and any path $\mathfrak{p}$, the result is true due to B.2. Suppose the result holds for round t1t-1, i.e., $\mathbb{P}^{t-1}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)$. For round $t$,
$\mathbb{P}^{t}(\mathfrak{p})$ $=\prod_{e\in\mathfrak{p}}\varphi^{t}(e)$
[…]
$=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\cdot\exp\Big(\eta\sum_{e\in\mathfrak{p}}\textsf{w}^{t-1}(e)\Big)\cdot\frac{\Gamma^{t-1}(d)}{\Gamma^{t-1}(s)},$
@@
$\Gamma^{t-1}(s)=\sum_{\mathfrak{p}^{\prime}}\prod_{e\in\mathfrak{p}^{\prime}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))=\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\cdot\exp\Big(\eta\sum_{e\in\mathfrak{p}^{\prime}}\textsf{w}^{t-1}(e)\Big)$
@@
Substituting the values gives $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. So, Algorithm 1 is a correct implementation of the Hedge algorithm.
@@
Regret Upper Bound and Time Complexity. Having shown that Algorithm 1 is equivalent to the Hedge algorithm, we now recall its regret bound from Cesa-Bianchi and Lugosi (2006, Theorem 2.2) which states that for learning rate $\eta$:
$\textsf{REG}\leq\frac{\log|\mathscr{P}|}{\eta}+\frac{\eta TM^{2}}{8}\,.$
As $|\mathscr{P}|\leq M^{m}$,
$\textsf{REG}\leq\frac{m\log M}{\eta}+\frac{\eta TM^{2}}{8}\,.$
Setting $\eta=\frac{1}{M}\sqrt{\frac{8m\log M}{T}}$, we get $\textsf{REG}\leq O(M\sqrt{mT\log M})$. In the full information setting, the running time bottleneck is computing Eq. 31. Thus, Algorithm 1 runs in $|E|=O(mM^{2})$ time per round.
B.3. Proof of Theorem 4.4
Regret Analysis. In the bandit setting, the bidder follows Algorithm 1 by substituting $\textsf{w}^{t}(e)$ with its unbiased estimator $\widehat{\textsf{w}}^{t}(e)$. Thus, in this setting, the updates is similar to that of the EXP3 algorithm (Auer et al., 2002). Using the standard analysis for EXP3 algorithm (see Lattimore and Szepesvári (2020, Chapter 11)), we obtain
@@
$\textsf{REG}=\frac{\log|\mathscr{P}|}{\eta}+\sum_{t=1}^{T}\mathbb{E}\left[\frac{1}{\eta}\log\left(\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})e^{\eta\widehat{\textsf{w}}^{t}(\mathfrak{p})}\right)-\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})\right],$
where $\widehat{\textsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{t}(e)$. Note that $\widehat{\textsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{t}(e)\leq\sum_{e\in\mathfrak{p}}\overline{\textsf{w}}(e)\leq M$ and $|\mathscr{P}|\leq M^{m}$. Note that if the natural importance-weighted estimator had been used instead, we would not be able to guarantee that $\widehat{\textsf{w}}^{t}(\mathfrak{p})\leq M$. For $\eta\leq\frac{1}{M}$, $\eta\widehat{\textsf{w}}^{t}(\mathfrak{p})\leq 1$. Hence,
REG $\leq\frac{m\log M}{\eta}+\eta\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\mathbb{E}[\widehat{\textsf{w}}^{t}(\mathfrak{p})^{2}]$
$\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}],$
@@
where in the first inequality, we used $e^{x}\leq 1+x+x^{2}$ for $x\leq 1$ and $\log(1+x)\leq x$ for $x\geq 0$. The second inequality follows from Cauchy-Schwarz. Observe that
$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}]=\sum_{e\in E}\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}]\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{e\in E}p^{t}(e)\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}],$
@@
$\sum_{e\in E}p^{t}(e)\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}]$ $=\sum_{e\in E}p^{t}(e)\Big[\overline{\textsf{w}}(e)^{2}(1-p^{t}(e))+\Big(\overline{\textsf{w}}(e)-\frac{\overline{\textsf{w}}(e)-\textsf{w}^{t}(e)}{p^{t}(e)}\Big)^{2}p^{t}(e)\Big]$
$=\sum_{e\in E}\textsf{w}^{t}(e)^{2}p^{t}(e)+(\overline{\textsf{w}}(e)-\textsf{w}^{t}(e))^{2}(1-p^{t}(e))\leq\sum_{e\in E}\overline{\textsf{w}}(e)^{2}\,.$
Observe that $\sum_{e}\overline{\mathsf{w}}(e)^{2}\leq|E|M^{2}\lesssim mM^{4}$. Hence,